OpenStudy (chris215):

Find d/dx integral from 1 to x^2 ln(t)dt

OpenStudy (trojanpoem):

\[\frac{ d }{ dx } \int\limits_{1}^{x^2} \ln t dt\]

OpenStudy (trojanpoem):

use integration by parts to find the integral of lnt with respect to t then put the limits : int lnt dt = tlnt - t from 1 to x^2 ln1 - 1 - [ x^2 lnx^2 - x^2] 0 - 1 - x^2 lnx^2 + x^2 -1 - x^2 lnx^2 + x^2 Now differentiate this : -2x lnx^2 - x^2 * 2x/x^2 + 2x

OpenStudy (chris215):

(-2 x log^2(x)-(x^2×2 x)/x^2+2 x) written liek this?

OpenStudy (irishboy123):

you can use Liebnitz - differentiation under the integral sign: if: \(\dfrac{dI}{dx} = \dfrac{d}{dx} \int\limits_{u(x)}^{v(x)} f(t) \; dt = \dfrac{d}{dx} \left[ F(v(x)) - F (u(x)) \right] \) then \(\dfrac{dI}{dx} = f(v) \dfrac{dv}{dx} - f(u) \dfrac{du}{dx} \)

OpenStudy (chris215):

ok I got 2xln(x^2)

OpenStudy (michele_laino):

that's right!

OpenStudy (chris215):

Thnaks guys!